# Toward Efficient and Robust Behavior Models for Multi-Agent Driving Simulation

### v3

@@ IV Experiments
Datasets: We evaluate our approach on two datasets: 1) the publicly available INTERACTION dataset [35], containing 38 25538\,255 trajectories ($4.15\text{\,}\%$ VRUs) across $11$ locations; and 2) an in-house dataset commercially licensed by DeepScenario, comprising 88 54088\,540 trajectories ($43.59\text{\,}\%$ VRUs) across $8$ locations. Both datasets were captured using drones, ensuring high-quality data, and include highly interactive scenarios such as merging lanes, unsignalized intersections, and roundabouts. We use $20\text{\,}\%$ of the recordings per location for validation and $30\text{\,}\%$ for testing. A more compact version of the DeepScenario dataset [11] is publicly available and can be used to obtain similar results.
Baseline Models: We compare against several lightweight baselines:
@@
    Constant Velocity (CV): A learning-free baseline where agents are assumed to continue moving forward at a constant velocity.
@@
    GraphAIRL [20]: A more sophisticated agent-centric model that leverages a vectorized scene representation [13] and attention-based interaction modeling. We evaluate two variants: 1) trained with $c=5$, as proposed in [20], and 2) trained with our proposed adaptive reward offset, defined in (3).
@@
Training Details: All models are trained using the AdamW optimizer [21] with a batch size of $1024$. Learning rates are set to $0.0002$ for the policy model and $0.0001$ for the discriminator, both decayed by a factor of $10$ during the final $30\text{\,}\%$ of training. For RL, we use advantage normalization. The models are trained for 10 00010\,000 epochs with a discount factor $\gamma=0.95$, GAE parameter $\lambda=0.95$, and target rewards $\tilde{r}_{\mathrm{target}}=19$ and $\tilde{r}_{\mathrm{target}}=11$ for INTERACTION and DeepScenario, respectively. For the policy model, the observation radius is set to $50\text{\,}\mathrm{m}$. Since the number of observed instances grows quadratically with the observation range, we reduce the discriminator’s observation radius to $30\text{\,}\mathrm{m}$, allowing it to focus primarily on the target agent and its direct surroundings.
In each epoch, scenarios with approximately $880$ agents are sampled from the training set. Validation scenarios contain roughly $1650$ agents, while test scenarios feature $5063$ and 19 41019\,410 agents for INTERACTION and DeepScenario, respectively. Agents are uniformly distributed across locations. Each scenario is simulated for $10\text{\,}\mathrm{s}$ at $5\text{\,}\mathrm{H}\mathrm{z}$. For final evaluation, we select the model with the lowest Root Mean Square Error (RMSE), normalized by (1Off-Track RateCollision Rate)(1-$\text${Off-Track Rate}-$\text${Collision Rate}), on the validation set, thus balancing prediction accuracy and robustness.
For our proposed model, we use $3$ Perceiver layers with a hidden dimension of $128$. We also evaluate a more compact variant with a single Perceiver layer and hidden dimension $64$, denoted Ours (small). All baseline models use a hidden dimension of $128$.
@@
Model Observation Num All (19 41019\,410 agents) Non-VRUs (10 94910\,949 agents) VRUs ($8461$ agents) Training Time [h]
Type Parameter RMSE [m] Off-Track [%] Collision [%] RMSE [m] Off-Track [%] Collision [%] RMSE [m] Collision [%] T4 16GB A100 80GB
Ground Truth - - 0.0 0.469 1.865 0.0 0.831 0.219 0.0 3.995 - -
CV - - 27.78 27.70 27.54 38.44 49.17 32.85 6.97 20.69 - -
[…]
Reported as mean $\pm$ std across 6 random seeds. $\dagger$}’ indicates that the baseline model was trained with our proposed target reward. AC = agent-centric. IC = instance-centric.
@@
Model RMSE [m] Off-Track [%] Collision [%]
Ground Truth 0.00 0.47 0.12
CV 23.17 28.23 27.74
[…]
Reported as mean $\pm$ std across 6 random seeds.
@@
Model RMSE [m] Off-Track [%] Collision [%]
Ground Truth 0.00 0.47 0.12
CV 23.17 28.23 27.74
[…]
Reported as mean $\pm$ std across 6 random seeds. Models are trained on DeepScenario dataset and evaluated on INTERACTION dataset.
@@ IV-A Results
We evaluate our proposed approach against a diverse set of baselines on two automated driving datasets. Table I presents the quantitative results on the DeepScenario dataset. Our proposed model achieves the lowest RMSE, off-track, and collision rates for both overall performance and the non-VRU subset. For VRUs, although our method shows slightly higher positional errors compared to some baselines, it is significantly more robust compared to all baselines.
Overall, our model outperforms all baselines on seven out of eight metrics, and even our small variant with only 5959K parameters surpasses all baselines on the same seven metrics. The only exception is VRU RMSE, where two baselines achieve marginally better positional accuracy. Remarkably, our approach exhibits greater robustness than ground truth agents, whose errors largely stem from annotation noise.
The learning-free CV baseline achieves reasonable positional accuracy for VRUs, as most do not abruptly change speed or direction, but performs poorly on other metrics, underscoring the need to model interactions. LateFusionMLP captures more realistic behaviors but still suffers from high off-track and collision rates, likely due to limited model capacity. Both GraphAIRL variants demonstrate high positional accuracy and robustness. Notably, training the model with our adaptive reward offset significantly improves its robustness while only slightly affecting its positional accuracy. Finally, the BC baseline achieves low RMSE but incurs extremely high off-track and collision rates due to the aforementioned covariate shift when being executed in simulation.
Table II presents the evaluation results on the INTERACTION dataset. Similarly, both our proposed models demonstrate high positional accuracy and robustness, outperforming all baselines on two of the three evaluated metrics.
Effects on Training Time: As shown in Table I, training time scales with model size and complexity. Both of our models achieve lower training times than all agent-centric baselines. Despite having more parameters, our instance-centric model consistently trains faster. The small variant (5959K parameters) reduces training time by up to $66.4\text{\,}\%$ compared to the agent-centric baselines, without compromising performance. Moreover, whereas agent-centric models typically require more computational power, our compact model achieves training times on a T4 GPU comparable to the baselines running on an A100 GPU.
@@
Effects on Inference Time: Figure 5 presents a quantitative comparison of single-step inference times across our models and baseline approaches. For the agent-centric models, inference time scales linearly with the number of agents, as observations must be encoded independently from each agent’s point of view. Nevertheless, the LateFusionMLP consistently exhibits low inference latency due to its compact MLP-based architecture. In contrast, our instance-centric approach decomposes inference into two stages: an initial simulation step, in which the entire map and all agents are encoded, and subsequent steps, in which the encoded map tokens can be reused, substantially reducing the inference time for later simulation steps. As a result, simulating $n$ timesteps incurs the latency of the initial step only once, followed by the latency of subsequent steps n1n-1 times. For scenarios with few agents, the initial forward pass of our proposed small model is slightly slower than that of GraphAIRL, since it encodes the entire map rather than only the regions visible to each agent. However, this overhead quickly diminishes as the number of simulation steps increases, yielding greater overall efficiency for longer simulations. Additionally, due to the symmetric scene context encoding, inference times in our instance-centric models stay nearly constant, regardless of how many agents are being simulated.
@@
Commonly, multiple simulations are run in parallel. To evaluate the sample throughput of the behavior model under parallel simulation, we introduce the metric Inference Steps per Second (ISPS). This metric quantifies the number of agents across all simulation environments that can be processed by the behavior model per second. Formally, it is defined as $\mathrm{ISPS}=H\times\sum_{k=1}^{N}N_{A}^{(k)}/\Delta T$, where $N_{A}^{(k)}$ denotes the number of agents in the $k$-th simulation environment, $H$ is the number of inference steps taken, and $\Delta T$ is the elapsed time in seconds. Figure 6 illustrates how the behavior model throughput scales with increasing numbers of parallel simulation environments as well as the distribution of simulated agents across $10^{3}$ randomly sampled scenarios. At smaller scales, all models benefit from parallelism, showing substantial gains in ISPS. Agent-centric models, however, saturate at around 27 00027\,000 and 36 00036\,000 ISPS, and could not be scaled beyond $10^{2}$ environments due to memory limits. In contrast, instance-centric models achieve higher throughout: both surpass agent-centric peaks with only ten parallel environments and continue scaling up to $10^{3}$. They reach peak throughput of about 88 00088\,000 and 358 000358\,000 ISPS, representing an up to $13.2$-fold improvement over the agent-centric baselines.
@@
Effects on Cross-Dataset Model Performance: Next, we evaluate the models performance in unseen traffic scenarios. For that, we train the models on the DeepScenario dataset and subsequentially evaluate the models on the INTERACTION dataset. As depicted in Table III, all models trained on the DeepScenario dataset, which predominantly contains scenarios from Germany, experience notable performance drops when evaluated on the INTERACTION dataset, which includes diverse scenarios from Germany, China, and the USA. This effect is particularly pronounced in the baseline models. In contrast, our proposed instance-centric models demonstrate greater robustness, maintaining relatively low collision and off-track rates while exhibiting smaller increases in RMSE compared to the baselines. We attribute this improved generalization to the viewpoint-invariant encoding of instances, which produces more transferable scene representations. By comparison, agent-centric approaches excel at modeling an individual agent’s local context but remain brittle to changes in viewpoint or environmental conditions, limiting their applicability to unseen scenarios. Notably, the collision rate of our models increases only marginally, highlighting their capacity to better generalize to diverse interactions with other traffic participants.
@@
Effects of Average Reward Target: Figure 7 presents the results of our ablation study on different choices for the average reward target $\tilde{r}_{\mathrm{target}}$, evaluated across values from $1$ to $29$ in increments of $2$. In general, increasing $\tilde{r}_{\mathrm{target}}$ leads to more robust behavior, with larger gains being achieved at lower values. Initially, the positional accuracy, measured by RMSE, increases significantly, demonstrating that the model benefits from the improved robustness. However, increasing $\tilde{r}_{\mathrm{target}}$ too much, leads to the positional error to rise again. For instance, this turning point can easily be seen for $\tilde{r}_{\mathrm{target}}=11$ on the DeepScenario dataset. This indicates a shift in the model’s behavior, where it starts to prioritize robustness over realism. Based on these results, we selected the values $11$ for DeepScenario and $19$ for INTERACTION, representing the optimal trade-off between prediction accuracy and robustness.
@@
Figure 7: Ablation on the validation set w. r. t. the average reward target defined in (3) for a) INTERACTION dataset and b) DeepScenario dataset. Results averaged over 4 seeds; whiskers indicate $\pm 1$ standard deviation.
@@
Example Situation: Figure 3 shows a highly interactive example situation featuring an all-way-stop intersection with unprotected left turns. The vehicles are controlled by our learned behavior model. Notably, the simulated and corresponding ground truth trajectories closely match, even over long horizons. The policy captures realistic driving dynamics, including similar speeds, accelerations, and subtle behaviors such as cooperative merging.

### v5

@@ IV Experiments
Datasets: Evaluation is conducted on two datasets: 1) the publicly available INTERACTION dataset [35], containing 38 25538\,255 trajectories ($4.15\text{\,}\%$ VRUs) across $11$ locations; and 2) an in-house dataset commercially licensed by DeepScenario, comprising 88 54088\,540 trajectories ($43.59\text{\,}\%$ VRUs) across $8$ locations. Both datasets were recorded using drones, ensuring high-quality data, and focus on interactive scenarios such as merging lanes, unsignalized intersections, and roundabouts. We use $20\text{\,}\%$ of the recordings per location for validation and $30\text{\,}\%$ for testing. A more compact version of the DeepScenario dataset [11] is publicly available and yields similar results.
Baselines: We compare against several compact baselines:
@@
    Constant Velocity (CV): A learning-free baseline that assumes agents continue forward at a constant velocity.
@@
    GraphAIRL [20]: A more sophisticated agent-centric model that uses a vectorized scene representation [13] and attention-based interaction modeling. We evaluate two variants: 1) trained with $c=5$, as done in [20], and 2) trained with our proposed adaptive reward offset (3).
@@
Training Details: All models are trained using the AdamW optimizer [21] with a batch size of $1024$. Learning rates are set to $0.0002$ for the policy model and $0.0001$ for the discriminator, both decayed by a factor of $10$ during the final $30\text{\,}\%$ of training. For RL, we use advantage normalization. The models are trained for 10 00010\,000 epochs with a discount factor $\gamma=0.95$, GAE parameter $\lambda=0.95$, and target rewards $\tilde{r}_{\mathrm{target}}=33$ and $\tilde{r}_{\mathrm{target}}=9$ for INTERACTION and DeepScenario, respectively. For the policy, the observation radius is set to $50\text{\,}\mathrm{m}$. Since the number of observed instances grows quadratically with the observation range, we reduce the discriminator’s observation radius to $30\text{\,}\mathrm{m}$, allowing it to focus primarily on the target agent and its direct surroundings.
In each epoch, scenarios with approximately $880$ agents are sampled from the training set. Validation scenarios contain roughly $1650$ agents, while test scenarios feature $5109$ and 19 63219\,632 agents for INTERACTION and DeepScenario, respectively. Agents are uniformly distributed across locations. Each scenario is simulated for $10\text{\,}\mathrm{s}$ at $5\text{\,}\mathrm{H}\mathrm{z}$. For final evaluation, we select the model that achieves the lowest aggregated score
$S(\theta)=\frac{\sqrt{\frac{1}{N_{a}}\sum_{n=1}^{N_{a}}\left(\mathrm{FDE}_{n}^{(\pi_{\theta})}\right)^{2}}}{\max(1-\mathrm{off}^{(\pi_{\theta})}-\mathrm{col}^{(\pi_{\theta})},\;\epsilon)},$ (8)
i. e., the root mean squared final displacement error normalized by (1Off-Track RateCollision Rate)(1-$\text${Off-Track Rate}-$\text${Collision Rate}). To ensure numerical stability, we set $\epsilon=10^{-6}$. The Final Displacement Error (FDE) is the Euclidean distance to the ground truth agent at the final simulation step. This score balances positional accuracy and robustness. During evaluation, terminated vehicles are not removed from simulation.
We use three Perceiver layers and a hidden dimension of $128$. Additionally, we evaluate a smaller variant with one Perceiver layer and a hidden dimension of $64$, denoted Ours (small). All baseline models use a hidden dimension of $128$.
@@
Model Observation Num FDE Collision w/ Off- Aggregated Training Time
Type Parameter All non-VRU VRU non-VRU VRU Route Track Score T4 16GB A100 80GB
Expert Demo - - $0.0$ $0.0$ $0.0$ $0.16$ $0.05$ $3.75$ $0.83$ $0.0$ - -
CV - - $16.73$ $29.18$ $\mathbf{3.99}$ $26.16$ $3.76$ $57.52$ $49.17$ $123.72$ - -
LateFusionMLP [9] AC 105105k $9.28_{\pm 0.14}$ $13.10_{\pm 0.23}$ $5.33_{\pm 0.20}$ $1.61_{\pm 0.23}$ $0.55_{\pm 0.09}$ $3.89_{\pm 0.56}$ $\underline{0.61}_{\pm 0.02}$ $14.59_{\pm 0.22}$ $86.85$ $40.75$
GraphAIRL [20] AC 146146k $7.86_{\pm 0.15}$ $10.78_{\pm 0.18}$ $4.83_{\pm 0.24}$ $1.09_{\pm 0.38}$ $0.79_{\pm 0.26}$ $3.96_{\pm 0.33}$ $0.70_{\pm 0.13}$ $12.58_{\pm 0.28}$ $125.94$ $45.95$
GraphAIRL [20] AC 146146k $7.79_{\pm 0.12}$ $10.84_{\pm 0.19}$ $4.60_{\pm 0.14}$ $0.67_{\pm 0.27}$ $0.42_{\pm 0.29}$ $4.16_{\pm 0.55}$ $\underline{0.61}_{\pm 0.03}$ $12.41_{\pm 0.24}$ $129.03$ $47.10$
BC IC 430430k $7.75_{\pm 0.53}$ $11.40_{\pm 0.97}$ $\underline{4.03}_{\pm 0.25}$ $14.71_{\pm 2.92}$ $2.90_{\pm 0.28}$ $16.72_{\pm 10.43}$ $10.51_{\pm 9.80}$ $17.35_{\pm 5.22}$ $12.67$ $3.07$
Ours (small) IC 6060k $\underline{7.59}_{\pm 0.23}$ $\underline{10.31}_{\pm 0.35}$ $4.74_{\pm 0.19}$ $\underline{0.66}_{\pm 0.33}$ $\underline{0.27}_{\pm 0.09}$ $\underline{3.85}_{\pm 0.70}$ $\mathbf{0.57}_{\pm 0.03}$ $\underline{12.01}_{\pm 0.34}$ $45.53$ $25.13$
Ours IC 430430k $\mathbf{7.09}_{\pm 0.23}$ $\mathbf{9.69}_{\pm 0.24}$ $4.43_{\pm 0.26}$ $\mathbf{0.38}_{\pm 0.13}$ $\mathbf{0.14}_{\pm 0.05}$ $\mathbf{3.27}_{\pm 0.23}$ $\mathbf{0.57}_{\pm 0.03}$ $\mathbf{11.29}_{\pm 0.35}$ $125.75$ $44.07$
Reported as mean $\pm$ std across $7$ random seeds. $\dagger$}’ indicates metric reported for the non-VRU subset. $\ddagger$}’ indicates baseline model trained with the proposed target reward. FDE in [m]. Collision, off-route and off-track rates in [%]. Training times in [h]. AC = agent-centric. IC = instance-centric.
@@
Model FDE Collision Off-Track Aggregated Training Time
Score T4 16GB A100 80GB
Expert Demo $0.0$ $0.0$ $0.51$ 0 - -
CV $17.44$ $22.59$ $30.65$ $49.55$ - -
LateFusionMLP [9] $10.12_{\pm 0.14}$ $0.57_{\pm 0.24}$ $\mathbf{0.30}_{\pm 0.02}$ $14.20_{\pm 0.30}$ $97.57$ $44.48$
GraphAIRL [20] $\underline{8.51}_{\pm 0.35}$ $1.38_{\pm 0.48}$ $0.36_{\pm 0.04}$ $\underline{12.49}_{\pm 0.48}$ $152.35$ $52.26$
GraphAIRL [20] $9.05_{\pm 0.47}$ $0.70_{\pm 0.25}$ $\mathbf{0.30}_{\pm 0.03}$ $12.94_{\pm 0.50}$ $157.19$ $55.14$
BC $10.28_{\pm 0.51}$ $13.98_{\pm 2.15}$ $9.65_{\pm 7.02}$ $19.09_{\pm 2.22}$ $2.37$ $0.68$
Ours (small) $8.78_{\pm 0.39}$ $\underline{0.40}_{\pm 0.22}$ $\underline{0.32}_{\pm 0.04}$ $12.56_{\pm 0.58}$ $42.47$ $24.17$
Ours $\mathbf{8.34}_{\pm 0.29}$ $\mathbf{0.38}_{\pm 0.18}$ $\mathbf{0.30}_{\pm 0.02}$ $\mathbf{12.08}_{\pm 0.26}$ $126.08$ $50.43$
Reported as mean $\pm$ std across $7$ random seeds.
@@
Model FDE Collision Off-Track Aggregated
Score
Expert Demo $0.0$ $0.0$ $0.51$ $0.0$
CV $17.44$ $22.59$ $30.65$ $49.55$
LateFusionMLP [9] $15.22_{\pm 1.34}$ $6.47_{\pm 1.97}$ $6.89_{\pm 2.85}$ $22.82_{\pm 3.21}$
GraphAIRL [20] $16.05_{\pm 1.14}$ $5.54_{\pm 2.57}$ $4.12_{\pm 2.13}$ $22.80_{\pm 2.62}$
GraphAIRL [20] $14.44_{\pm 0.49}$ $3.24_{\pm 0.84}$ $3.89_{\pm 3.06}$ $20.43_{\pm 1.44}$
BC $15.62_{\pm 1.31}$ $19.69_{\pm 1.46}$ $13.95_{\pm 3.65}$ $30.25_{\pm 2.15}$
Ours (small) $\underline{14.26}_{\pm 1.48}$ $\underline{0.81}_{\pm 0.28}$ $\underline{1.49}_{\pm 0.79}$ $\underline{19.01}_{\pm 1.72}$
Ours $\mathbf{12.62}_{\pm 0.69}$ $\mathbf{0.51}_{\pm 0.16}$ $\mathbf{1.11}_{\pm 0.16}$ $\mathbf{16.82}_{\pm 0.71}$
Reported as mean $\pm$ std across $7$ random seeds. Models are trained on DeepScenario dataset and evaluated on INTERACTION dataset.
@@ IV-A Results
We evaluate our proposed approach against a diverse set of baselines on two automated driving datasets. Table I presents the quantitative results on the DeepScenario dataset. The proposed model achieves the lowest FDE for the overall and non-VRU subset, as well as the most robust behavior in terms of collision, off-route and off-track rates. Overall, our model outperforms all baselines on seven out of eight metrics, and even our small variant with only 6060k parameters surpasses all baselines on the same seven metrics. The only exception is VRU FDE, where two baselines achieve marginally better positional accuracy. Notably, our approach achieves lower off-route and off-track rates than ground truth agents, whose errors largely stem from annotation noise.
The learning-free CV baseline achieves reasonable positional accuracy for VRUs, as most do not abruptly change speed or direction, but performs poorly on other metrics, underscoring the need to model interactions. LateFusionMLP captures more realistic behaviors but still suffers from low robustness and comparably high FDE, likely due to limited model capacity. Both GraphAIRL variants demonstrate high positional accuracy and robustness. Notably, training the model with our adaptive reward offset significantly improves its robustness and yields a small gain in positional accuracy. Finally, the BC baseline achieves low FDE but incurs extremely high off-track and collision rates due to the aforementioned covariate shift when executed in simulation.
Table II presents the evaluation results on the INTERACTION dataset. Similarly, the proposed model achieves the best performance on all four evaluation metrics. Its compact variant Ours (small) remains competitive, exhibiting comparable FDE and outperforming all baselines in robustness.
Effects on Training Time: As shown in Tables I and II, training time scales with model size and architectural complexity. Despite having more parameters, Ours outperforms both GraphAIRL baselines in terms of training times. Ours (small) consistently trains faster than all agent-centric baselines, reducing training time by up to $64.71\text{\,}\%$ on DeepScenario and $72.98\text{\,}\%$ on INTERACTION without compromising performance. Moreover, while agent-centric models typically require more computational power, our compact model variant achieves training times on a T4 $16\text{\,}\mathrm{GB}\text{/}$ GPU comparable to the baselines trained on an A100 $80\text{\,}\mathrm{GB}\text{/}$ GPU.
@@
Effects on Inference Time: Figure 5 presents a quantitative comparison of single-step inference times across our models and baseline approaches. For the agent-centric models, inference time scales linearly with the number of agents, as observations must be encoded separately from each agent’s point of view. Nevertheless, the LateFusionMLP consistently exhibits low inference latency due to its compact MLP-based architecture. In contrast, our instance-centric approach decomposes inference into two stages: an initial simulation step, in which the entire map and all agents are encoded, and subsequent steps, in which the encoded map tokens are reused, substantially reducing the inference time for later simulation steps. As a result, simulating $n$ steps incurs the latency of the initial step only once, followed by the latency of subsequent steps n1n-1 times. For scenarios with few agents, the initial forward pass of our proposed small model is slightly slower than that of GraphAIRL, since it encodes the entire map rather than only the regions visible to each agent. However, this overhead quickly diminishes as the number of simulation steps increases, yielding greater overall efficiency for longer simulations. Additionally, due to the symmetric scene context encoding, inference times in our instance-centric models stay nearly constant, regardless of the number of simulated agents.
@@
Commonly, multiple simulations are run in parallel. To evaluate the sample throughput of the behavior model under parallel simulation, we introduce the metric Inference Steps per Second (ISPS). This metric quantifies the number of agents across all simulation environments that can be processed by the behavior model per second. Formally, it is defined as $\mathrm{ISPS}=H\times\sum_{k=1}^{N}N_{A}^{(k)}/\Delta T$, where $N_{A}^{(k)}$ denotes the number of agents in the $k$-th simulation environment, $H$ is the number of inference steps taken, and $\Delta T$ is the elapsed time in seconds. Figure 6 illustrates how the behavior model throughput scales with increasing numbers of parallel simulation environments as well as the distribution of simulated agents across $10^{4}$ randomly sampled scenarios. At smaller scales, all models benefit from parallelism, showing substantial gains in ISPS. Agent-centric models saturate early at around 28.728.7k and 60.360.3k ISPS, and cannot be scaled beyond $500$ environments due to memory limits. In contrast, instance-centric models achieve higher throughout: Ours (small) and Ours exceed both agent-centric models with only five and twenty parallel environments, respectively, and continue scaling up to $5000$ parallel environments. They reach peak throughput of about 93.493.4k and 340.7340.7k ISPS, representing an up to $11.9$-fold improvement over the agent-centric baselines.
@@
Effects on Cross-Dataset Performance: Next, we evaluate the models’ performance in unseen traffic scenarios. For that, the models are trained on the DeepScenario dataset and subsequently evaluated on the INTERACTION dataset. As depicted in Table III, all models trained on the DeepScenario dataset, which predominantly contains scenarios from Germany, experience notable performance drops when evaluated on the INTERACTION dataset, which includes diverse scenarios from Germany, China, and the USA. This effect is particularly pronounced in the baseline models. In contrast, our proposed instance-centric model demonstrates greater robustness, maintaining relatively low collision and off-track rates while exhibiting smaller increases in FDE compared to the baselines. We attribute this improved generalization to the viewpoint-invariant encoding of instances, which produces more transferable scene representations. By comparison, agent-centric approaches excel at modeling an individual agent’s local context but remain brittle to changes in viewpoint or environmental conditions, limiting applicability to unseen scenarios. Notably, the collision rate of our models increases only marginally, highlighting their capacity to better generalize to diverse interactions with other agents.
@@
Effects of Average Reward Target: Figure 7 presents the results of our ablation study on different choices for the average reward target $\tilde{r}_{\mathrm{target}}$, evaluated across values from $1$ to $39$ in increments of $2$. In general, increasing $\tilde{r}_{\mathrm{target}}$ leads to more robust behavior, with larger gains being achieved at lower values. Initially, the positional accuracy, measured by FDE, increases significantly, demonstrating that the model benefits from the improved robustness. However, increasing $\tilde{r}_{\mathrm{target}}$ too much, leads to the positional error to rise again. For instance, this turning point can easily be seen for $\tilde{r}_{\mathrm{target}}=9$ on the DeepScenario dataset. This indicates a shift in the model’s behavior, where it starts to prioritize robustness over realism. Based on these results, we selected the values $9$ for DeepScenario and $33$ for INTERACTION, representing the optimal trade-off between positional accuracy and robustness.
@@
Figure 7: Ablation on the validation set w. r. t. the average reward target defined in (3) for a) INTERACTION dataset and b) DeepScenario dataset. Collision and off-track rates are calculated on the non-VRU subset. Results averaged over 55 seeds. Whiskers indicate $\pm 1$ standard deviation.
@@
Example Scenario: Figure 3 shows a highly interactive example scenario featuring an all-way-stop intersection. All agents are controlled by our learned behavior model. Notably, the simulated and corresponding ground truth trajectories closely match, even over long horizons. The policy captures realistic driving dynamics, including similar speeds, accelerations, and subtle behaviors such as cooperative merging.